If 3 is added to a number and this sum is doubled, the result is 2 more than the number. What is the equation. and what does it mean when it says "sum is doubled?"
Answer by checkley77(12844)

2(x+3)=x+2
2x+6=x+2
2x-x=2-6
x=-4 ans.
Proof:
2(-4+3)=-4+2
2*-1=-2
-2=-2